OpenStudy (anonymous):

Condence The Expression 2ln x-3 lny

OpenStudy (anonymous):

ln(x^2-y^3)

OpenStudy (anonymous):

How Did You Get That Answer ?

OpenStudy (anonymous):

according to the property of log function

OpenStudy (anonymous):

2logx=logx^2

OpenStudy (anonymous):

logx+logy=log(xy), logx-logy=log(x/y)